OpenStudy (anonymous):

i will give medal write an equation in point slope form for the line through the given points (-4, 6) m = 3/4]

OpenStudy (jidnyesha):

\[y=\frac{ 3 }{ 4 }x+9\]

OpenStudy (jidnyesha):

y=mx+c y=6 and x=-4 so \[6=\frac{ 3 }{ 4 }*(-4)+c\] \[6=-3+c\] c=6+3 c=9 so, \[y=\frac{ 3 }{ 4 }x+9\]

OpenStudy (anonymous):

point slope form : y - y1 = m(x - x1) m = 3/4 (-4,6) x1 = -4 and y1 = 6 now we sub y - 6 = 3/4(x - (-4) y - 6 = 3/4(x + 4) ===> point slope form

OpenStudy (pratyush5):

@Jidnyesha ut answer is correct but that is slope intercept from . @kelliegirl33 is right.

OpenStudy (anonymous):

@kelliegirl33 so answer is y - 6 = 3/4(x + 4)

OpenStudy (anonymous):

yes

OpenStudy (jidnyesha):

i dont actually know the difference between slope intercept and point intercept

OpenStudy (anonymous):

it can get confusing

OpenStudy (anonymous):

y = mx + b is slope intercept form y - y1 = m(x - x1) is point slope form

OpenStudy (jidnyesha):

when i solve the equation u said was right i get the equation i posted

OpenStudy (anonymous):

can someone help with this question

OpenStudy (pratyush5):

Actually slope intercept form is derived from slope point form.

OpenStudy (jidnyesha):

ohk when v were taught about the graphs nothing was mentioned about point intercept

OpenStudy (anonymous):

Let the time be x, and height y. Substitute those values in y=mx + c or slope formula, then solve simultaneously.

OpenStudy (anonymous):

take 2 points from your chart...(2,14) and (4,23) find slope slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (23 - 14) / (4 - 2) slope(m) = 9/2 y = mx + b slope(m) = 9/2 (2,14) x = 2 and y = 14 now we sub 14 = 9/2(2) + b 14 = 18/2 + b 14 = 9 + b 14 - 9 = b 5 = b your slope intercept equation is : y = 9/2x + 5 do you understand ? Do you have any questions ?

OpenStudy (anonymous):

yep. nope thanks